OpenStudy (anonymous):

any one know how to solve this problem? x^2 + 2x = 0

OpenStudy (zpupster):

pull out an x x(x+2)=0 then set the 2 terms equal to 0 x=0 and x+2 = 0 subtract 2 from both sides x+2-2 = 0-2 x=-2

OpenStudy (anonymous):

i have to factor the left side of the equation then use the zero product property to find all real number solutions

OpenStudy (zpupster):

sounds good

OpenStudy (anonymous):

how do you do that?

OpenStudy (anonymous):

help please

OpenStudy (agent0smith):

@crissy126 that is exactly what zpupster just did.

OpenStudy (anonymous):

ok thx